# Optimal upper bounds for anti-canonical volumes of singular toric Fano varieties

### v1

@@ 3. Inequalities for barycentric coordinates
Fix two integers $d\geq 2$ and $q\geq 1$. Let $S$ be a $d$-dimensional $\frac{1}{q}$-lc Fano simplex. Suppose $\beta_{1}\geq\beta_{2}\geq...\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}\in\operatorname{int}(S)$ with respect to the vertices $v_{1},v_{2},...,v_{d+1}$. In this section, we will give lower bounds for $\Pi_{i=1}^{d}\beta_{i}$ and $\Pi_{i=1}^{d+1}\beta_{i}$, which are the main ingredients for the proof of the main results.
By doing the similar operation as in the proof of [2, Theorem 1.1], we can get generalized Product-Sum inequalities for barycentric coordinates.
@@
From the proof we can see that the conclusions also hold for a $d$-dimensional latttice simplex $S$ in ${\mathbf{N}}_{\mathbb{R}}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$.
As indicated by Theorem 3.1, in order to find lower bounds for $\Pi_{i=1}^{d}\beta_{i}$ and $\Pi_{i=1}^{d+1}\beta_{i}$, we need to consider the set of $d+1$-tuples $(x_{1},x_{2},\cdots,x_{d+1})\in\mathbb{R}^{d+1}$ such that the following conditions hold:
@@
We denote this set by $X(d,q)$. Then $(\beta_{1},\cdots,\beta_{d+1})\in X(d,q)$ by Theorem 3.1. For convenience, let $PS[q](x)_{t}(t\in\{1,2,\cdots,d\})$ be the $t$-th Product-Sum inequality $\Pi_{i=1}^{t}x_{i}\leq q^{t}\Sigma_{j=t+1}^{d+1}x_{j}$ with respect to the positive integer $q$ and the $d+1$-tuple $x=(x_{1},x_{2},\cdots,x_{d+1})\in\mathbb{R}^{d+1}$. By definition, for any $x\in X(d,q)$, $PS[q](x)_{t}(t\in\{1,2,\cdots,d\})$ holds. And $x\in X(d,q)$ also satisfies the following basic properties (a generalized version of [3, Lemma 4.1]):
@@
    If the equality in $PS[q](x)_{t}$ holds for some $t\in\{1,\cdots,d\}$, then $x_{t}>x_{t+1}$;
@@
    Fix $k\in\{1,\cdots,d\}$, if the equality in $PS[q](x)_{t}$ holds for any $t\in\{1,\cdots,k\}$, then $x_{t}=\frac{q}{1+u_{t,q}}$ for each $t\in\{1,\cdots,k\}$.
@@
$(1)$ If $x_{1}=1$, then (3.2) and (3.3) imply $x_{i}=0$ for all $i\in\{2,\cdots,d+1\}$, which contradicts to $PS[q](x)_{1}$. If $x_{d+1}=0$, then (3.3) and $PS[q](x)_{d}$ imply that $x_{d}=0$. Inductively we can apply (3.3) and (3.4) to deduce that $x_{i}=0$ for all $i\in\{1,\cdots,d+1\}$, which is a contradiction.
@@
$(3)$ For $t=1$, $x_{1}=q\Sigma_{j=2}^{d+1}x_{j}$ implies $x_{1}=q(1-x_{1})$, and hence $x_{1}=\frac{q}{1+u_{1,q}}$. Assume $x_{i}=\frac{q}{1+u_{i,q}}(1\leq i\leq t)$ for some $11\leq t\leq k-$ 1. Then $\Pi_{i=1}^{t+1}x_{i}=q^{t+1}\Sigma_{j=t+2}^{d+1}x_{j}$ implies $(\Pi_{i=1}^{t}\frac{q}{1+u_{i,q}})x_{t+1}=q^{t+1}(1-\Sigma_{i=1}^{t}\frac{q}{1+u_{i,q}}-x_{t+1})$. Applying (1.1) and (1.2) with $p=t+1$, we can get $x_{t+1}=\frac{q}{1+u_{t+1,q}}.$ Then we can conclude the proof by induction. ∎
Now we come to study the lower bounds of $\Pi_{i=1}^{d}\beta_{i}$ and $\Pi_{i=1}^{d+1}\beta_{i}$.
@@
Let $f_{d+1}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d+1}(x)=\Pi_{i=1}^{d+1}x_{i}$. Let $b$ be the $d+1$-tuple $(b_{1},\cdots,b_{d+1})\in X(d,q)$ such that $f_{d+1}(b)=\min\{f_{d+1}(x)|x\in X(d,q)\}$. Then there exists an integer $k\in\{0,1,\cdots,d\}$ such that
@@
If the conclusion does not hold, then we may find $1\leq i_{1}<i_{2}\leq k$ such that $b_{i_{1}-1}>b_{i_{1}}=\cdots=b_{i_{2}}>b_{i_{2}+1}$. Take a real number $\delta>0$ and set $b^{\prime}:=(b_{1},\cdots,b_{i_{1}}+\delta,\cdots,b_{i_{2}}-\delta,\cdots,b_{d+1})$, i.e., $b^{\prime}$ is attained from $b$ by adding $\delta$ to $b_{i_{1}}$ and subtracting $b_{i_{2}}$ by $\delta$. Then $b^{\prime}$ satisfies condition (3.2). And condition (3.3) holds for $b^{\prime}$ if $\delta$ is sufficiently small. Furthermore, $b^{\prime}$ aslo satisfies condition (3.4) for a sufficiently small $\delta$: in fact, for $1\leq l<i_{1}$, $PS[q](b^{\prime})_{l}$ holds because it is the same as $PS[q](b)_{l}$; for $i_{1}\leq l<i_{2}$, note that by Lemma 3.3 $(2)$, $PS[q](b)_{l}$ is strict, so $PS[q](b^{\prime})_{l}$ holds if we take $\delta>0$ small enough; for $i_{2}\leq l\leq d$, the left product part of $PS[q](b^{\prime})_{l}$ is strict less than that of $PS[q](b)_{l}$ by the fact that $(b_{i_{1}}+\delta)(b_{i_{2}}-\delta)<b_{i_{1}}b_{i_{2}}$, and the right sum parts of them are the same, so $PS[q](b^{\prime})_{l}$ holds. Then for a sufficiently small real number $\delta>0$, we have $b^{\prime}\in X(d,q)$. But $(b_{i_{1}}+\delta)(b_{i_{2}}-\delta)<b_{i_{1}}b_{i_{2}}$ implies $f_{d+1}(b^{\prime})<f_{d+1}(b)$, which contradicts the choice of $b$. So we finish the proof of the claim. ∎
@@
If $k\geq 2$, then the equality in $PS[q](b)_{l}$ holds for any $11\leq l\leq k-$ 1.
@@
Suppose $PS[q](b)_{l}$ is strict for some $11\leq l\leq k-$ 1. Set $b^{*}:=(b_{1},\cdots,b_{l}+\delta,b_{l+1}-\delta,\cdots,b_{d+1})$ for some $\delta>0$. Then $b^{*}$ satisfies condition (3.2). By claim 3.5 and (3.5), we have $b_{l-1}>b_{l}$ and $b_{l+1}>b_{l+2}$, which imply condition (3.3) holds for $b^{*}$ if $\delta>0$ is small enough. By comparing $PS[q](b^{*})_{t}$ and $PS[q](b)_{t}$ for each $1\leq t\leq d$, we can see that $PS[q](b^{*})_{t}$ holds if $\delta>0$ is small enough. Then for a sufficiently small real number $\delta>0$, we have $b^{*}\in X(d,q)$. But $f_{d+1}(b^{*})<f_{d+1}(b)$ by construction, contradicting to the choice of $b$. ∎
To finish the proof, we need to show that the equality in $PS[q](b)_{k}$ holds. Suppose to the contrary that $\Pi_{i=1}^{k}b_{i}<q^{t}\Sigma_{j=k+1}^{d+1}b_{j}$. Set $b(\delta):=(b_{1},\cdots,b_{k}-(d+1-k)\delta,b_{k+1}+\delta,\cdots,b_{d+1}+\delta)$. Note that $PS[q](b)_{l}$ is strict for each $k\leq l\leq d$ by Lemma 3.3 $(2)$ and the assumptions. Arguing similarly as in the proof of the above two claims, we can have $b(\delta)\in X(d,q)$ as $\delta$ varies in a small open neighbourhood of 00. Denote $g(\delta):=\frac{f_{d+1}(b(\delta))}{\Pi_{i=1}^{k-1}b_{i}}=(b_{d+1}+\delta)^{d+1-k}\cdot(b_{k}-(d+1-k)\delta)$. Then $g(\delta)$ has a local minimal value at $\delta=0$. So the derivative value of $g(\delta)$ at 0 must be 00, which implies $b_{k}=b_{d+1}$, a contradiction.
@@
Let $f_{d}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d}(x)=\Pi_{i=1}^{d}x_{i}$. Let $a$ be the $d+1$-tuple $(a_{1},\cdots,a_{d+1})\in X(d,q)$ such that $f_{d}(a)=\min\{f_{d}(x)|x\in X(d,q)\}$. Then there exists an integer $k\in\{0,1,\cdots,d-1\}$ such that
@@
First we claim that $a_{d}=a_{d+1}$ always holds. Suppose $a_{d}>a_{d+1}$. Then one can take a small $\delta>0$ such that $a_{d}-\delta>a_{d+1}+\delta$. Set $a^{*}=(a_{1},\cdots,a_{d-1},a_{d}-\delta,a_{d+1}+\delta)$. We have $a^{*}\in X(d,q)$ and $f_{d}(a^{*})<f_{d}(a)$, contradicting to the choice of $a$.
@@
and the equality in $\text{the equality in }PS[q](a)_{l}\text{ holds for }1\leq l\leq k-1.$ So it suffices to prove that the equality in $PS[q](a)_{k}$ holds. Suppose to the contrary that $PS[q](a)_{k}$ is strict. For a real number $\delta$, set $a(\delta):=(a_{1},\cdots,a_{k}-(d+1-k)\delta,a_{k+1}+\delta,\cdots,a_{d+1}+\delta)$ and denote $h(\delta):=\frac{f_{d}(a(\delta))}{\Pi_{i=1}^{k-1}a_{i}}=(a_{d+1}+\delta)^{d-k}\cdot(a_{k}-(d+1-k)\delta)$. Arguing similarly as before, we can have $a(\delta)\in X(d,q)$ as $\delta$ varies in a small open neighbourhood of 00. Consider the derivative of $h(\delta)$
@@
By construction $h(\delta)$ has a local minimal value at $\delta=0$, we must have $h^{\prime}(0)=0$, which implies $a_{d+1}(d+1-k)=a_{k}(d-k)$. Consider the double derivative value of $h(\delta)$ at $\delta=0$: $h^{\prime\prime}(0)=$
$\begin{cases}(d-k)(d-1-k)a_{d+1}^{d-2-k}a_{k}-2(d-k)(d+1-k)a_{d+1}^{d-1-k}&\text{if }k<d-1;\\ -4&\text{if }k=d-1.\end{cases}$
@@
for each $l\in\{1,\cdots,d+1\}$. It is easy to see that for each $l\in\{1,\cdots,d+1\}$, $y[q](l)\in X(d,q)$.
@@
Keep the same notation as in the beginning of this section. Let $f_{d+1}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d+1}(x)=\Pi_{i=1}^{d+1}x_{i}$. Then
$\Pi_{i=1}^{d+1}\beta_{i}\geq f_{d+1}(y[q](d+1))=\frac{q^{d+2}}{u_{d+1,q}^{2}}.$
@@
One can check directly that equality (3.6) holds. So we may assume that $d\geq 3$.
In case $q=1$, the equality (3.6) follows from [3, Theorem 4.3 $(c)$]. So in the next discussion, we always assume that $q\geq 2$.
@@
Keep the same notation as in the beginning of this section. Let $f_{d}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d}(x)=\Pi_{i=1}^{d}x_{i}$. Then
@@
except for the case $d=2,q=1$. In the exceptional case $d=2,q=1$, $\Pi_{i=1}^{d}\beta_{i}\geq f_{2}(y[q](1))=\frac{1}{9}.$

### v2

@@ 3. Inequalities for barycentric coordinates
Fix two integers $d\geq 2$ and $q\geq 1$. Let $S$ be a $d$-dimensional $\frac{1}{q}$-lc Fano simplex. Suppose $\beta_{1}\geq\beta_{2}\geq...\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}\in\operatorname{int}(S)$ with respect to the vertices $v_{1},v_{2},...,v_{d+1}$. In this section, we will give lower bounds for $\Pi_{i=1}^{d+1}\beta_{i}$ and $\Pi_{i=1}^{d}\beta_{i}$, which are the main ingredients for the proof of the main results.
By generalizing the proof of [2, Theorem 1.1] to the situation considered here, we can get generalized Product-Sum inequalities for barycentric coordinates.
@@
From the proof we can see that the conclusions also hold for a $d$-dimensional latttice simplex $S$ in $\mathbb{R}^{d}$ such that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$.
As indicated by Theorem 3.1, in order to find lower bounds for $\Pi_{i=1}^{d+1}\beta_{i}$ and $\Pi_{i=1}^{d}\beta_{i}$, we need to consider the set of $(d+1)$-tuples $(x_{1},x_{2},\cdots,x_{d+1})\in\mathbb{R}^{d+1}$ such that the following conditions hold:
@@
We denote this set by $X(d,q)$. Then $(\beta_{1},\cdots,\beta_{d+1})\in X(d,q)$ by Theorem 3.1. For convenience, let $\operatorname{PS[q]}(x)_{t}(t\in\{1,2,\cdots,d\})$ be the $t$-th Product-Sum inequality $\Pi_{i=1}^{t}x_{i}\leq q^{t}\Sigma_{j=t+1}^{d+1}x_{j}$ with respect to the positive integer $q$ and the $(d+1)$-tuple $x=(x_{1},x_{2},\cdots,x_{d+1})\in\mathbb{R}^{d+1}$. By definition, for any $x\in X(d,q)$, $\operatorname{PS[q]}(x)_{t}(t\in\{1,2,\cdots,d\})$ holds. And $x\in X(d,q)$ also satisfies the following basic properties (a generalized version of [3, Lemma 4.1]):
@@
    If the equality in $\operatorname{PS[q]}(x)_{t}$ holds for some $t\in\{1,\cdots,d\}$, then $x_{t}>x_{t+1}$;
@@
    Fix $k\in\{1,\cdots,d\}$, if the equality in $\operatorname{PS[q]}(x)_{t}$ holds for any $t\in\{1,\cdots,k\}$, then $x_{t}=\frac{q}{1+u_{t,q}}$ for each $t\in\{1,\cdots,k\}$.
@@
$(1)$ If $x_{1}=1$, then (3.2) and (3.3) imply $x_{i}=0$ for all $i\in\{2,\cdots,d+1\}$, which contradicts $\operatorname{PS[q]}(x)_{1}$. If $x_{d+1}=0$, then (3.3) and $\operatorname{PS[q]}(x)_{d}$ imply that $x_{d}=0$. Inductively we can apply (3.3) and (3.4) to deduce that $x_{i}=0$ for all $i\in\{1,\cdots,d+1\}$, which is a contradiction.
@@
$(3)$ For $t=1$, $x_{1}=q\Sigma_{j=2}^{d+1}x_{j}$ implies $x_{1}=q(1-x_{1})$, and hence $x_{1}=\frac{q}{1+u_{1,q}}$. Assume $k\geq 2$ and for some $11\leq t\leq k-$ 1, we have $x_{i}=\frac{q}{1+u_{i,q}}$ for each $i\in\{1,\cdots,t\}$. Then $\Pi_{i=1}^{t+1}x_{i}=q^{t+1}\Sigma_{j=t+2}^{d+1}x_{j}$ implies $(\Pi_{i=1}^{t}\frac{q}{1+u_{i,q}})x_{t+1}=q^{t+1}(1-\Sigma_{i=1}^{t}\frac{q}{1+u_{i,q}}-x_{t+1})$. Applying (1.1) and (1.2) with $p=t+1$, we can get $x_{t+1}=\frac{q}{1+u_{t+1,q}}.$ So we conclude the proof by induction. ∎
Now we come to study the lower bounds of $\Pi_{i=1}^{d+1}\beta_{i}$ and $\Pi_{i=1}^{d}\beta_{i}$.
@@
Let $f_{d+1}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d+1}(x)=\Pi_{i=1}^{d+1}x_{i}$. Let $b$ be the $(d+1)$-tuple $(b_{1},\cdots,b_{d+1})\in X(d,q)$ such that $f_{d+1}(b)=\min\{f_{d+1}(x)|x\in X(d,q)\}$. Then there exists an integer $k\in\{0,1,\cdots,d\}$ such that
@@
If the conclusion does not hold, then we may find $1\leq i_{1}<i_{2}\leq k$ such that $b_{i_{1}-1}>b_{i_{1}}=\cdots=b_{i_{2}}>b_{i_{2}+1}$. Take a real number $\delta>0$ and set $b^{\prime}:=(b_{1},\cdots,b_{i_{1}}+\delta,\cdots,b_{i_{2}}-\delta,\cdots,b_{d+1})$, i.e., $b^{\prime}$ is attained from $b$ by adding $\delta$ to $b_{i_{1}}$ and subtracting $b_{i_{2}}$ by $\delta$. Then $b^{\prime}$ satisfies condition (3.2). And condition (3.3) holds for $b^{\prime}$ if $\delta$ is sufficiently small. Furthermore, $b^{\prime}$ aslo satisfies condition (3.4) for a sufficiently small $\delta$: in fact, for $1\leq l<i_{1}$, $\operatorname{PS[q]}(b^{\prime})_{l}$ holds because it is the same as $\operatorname{PS[q]}(b)_{l}$; for $i_{1}\leq l<i_{2}$, note that by Lemma 3.3 $(2)$, $\operatorname{PS[q]}(b)_{l}$ is strict, so $\operatorname{PS[q]}(b^{\prime})_{l}$ holds if we take $\delta>0$ small enough; for $i_{2}\leq l\leq d$, the left product part of $\operatorname{PS[q]}(b^{\prime})_{l}$ is strict less than that of $\operatorname{PS[q]}(b)_{l}$ by the fact that $(b_{i_{1}}+\delta)(b_{i_{2}}-\delta)<b_{i_{1}}b_{i_{2}}$, and the right sum parts of them are the same, so $\operatorname{PS[q]}(b^{\prime})_{l}$ holds. Hence for a sufficiently small real number $\delta>0$, we have $b^{\prime}\in X(d,q)$. But $(b_{i_{1}}+\delta)(b_{i_{2}}-\delta)<b_{i_{1}}b_{i_{2}}$ implies $f_{d+1}(b^{\prime})<f_{d+1}(b)$, which contradicts the choice of $b$. So we finish the proof of the claim. ∎
@@
If $k\geq 2$, then the equality in $\operatorname{PS[q]}(b)_{l}$ holds for any $11\leq l\leq k-$ 1.
@@
Suppose $\operatorname{PS[q]}(b)_{l}$ is strict for some $11\leq l\leq k-$ 1. Set $b^{*}:=(b_{1},\cdots,b_{l}+\delta,b_{l+1}-\delta,\cdots,b_{d+1})$ for some $\delta>0$. Then $b^{*}$ satisfies condition (3.2). By claim 3.5 and (3.5), we have $b_{l-1}>b_{l}$ and $b_{l+1}>b_{l+2}$, which imply condition (3.3) holds for $b^{*}$ if $\delta>0$ is small enough. By comparing $\operatorname{PS[q]}(b^{*})_{t}$ and $\operatorname{PS[q]}(b)_{t}$ for each $1\leq t\leq d$, we can see that $\operatorname{PS[q]}(b^{*})_{t}$ holds if $\delta>0$ is small enough. Hence for a sufficiently small real number $\delta>0$, we have $b^{*}\in X(d,q)$. But $f_{d+1}(b^{*})<f_{d+1}(b)$ by construction, contradicting the choice of $b$. ∎
To finish the proof, we need to show that the equality in $\operatorname{PS[q]}(b)_{k}$ holds. Suppose to the contrary that $\Pi_{i=1}^{k}b_{i}<q^{t}\Sigma_{j=k+1}^{d+1}b_{j}$. For a real number $\delta$, set
$b(\delta):=(b_{1},\cdots,b_{k}-(d+1-k)\delta,b_{k+1}+\delta,\cdots,b_{d+1}+\delta)$
and
$g(\delta):=\frac{f_{d+1}(b(\delta))}{\Pi_{i=1}^{k-1}b_{i}}=(b_{d+1}+\delta)^{d+1-k}\cdot(b_{k}-(d+1-k)\delta).$
Note that $\operatorname{PS[q]}(b)_{l}$ is strict for each $k\leq l\leq d$ by Lemma 3.3 $(2)$ and the assumptions. Arguing similarly as in the proof of the above two claims, we can have $b(\delta)\in X(d,q)$ as $\delta$ varies in a small open neighbourhood of 00. By construction $g(\delta)$ has a local minimal value at $\delta=0$. So the derivative value of $g(\delta)$ at 0 must be 00, which implies $b_{k}=b_{d+1}$, a contradiction.
@@
Let $f_{d}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d}(x)=\Pi_{i=1}^{d}x_{i}$. Let $a$ be the $(d+1)$-tuple $(a_{1},\cdots,a_{d+1})\in X(d,q)$ such that $f_{d}(a)=\min\{f_{d}(x)|x\in X(d,q)\}$. Then there exists an integer $k\in\{0,1,\cdots,d-1\}$ such that
@@
First we claim that $a_{d}=a_{d+1}$ always holds. Suppose $a_{d}>a_{d+1}$. Then one can take a small $\delta>0$ such that $a_{d}-\delta>a_{d+1}+\delta$. Set $a^{*}=(a_{1},\cdots,a_{d-1},a_{d}-\delta,a_{d+1}+\delta)$. We have $a^{*}\in X(d,q)$ and $f_{d}(a^{*})<f_{d}(a)$, contradicting the choice of $a$.
@@
and $\text{the equality in }\operatorname{PS[q]}(a)_{l}\text{ holds for }1\leq l\leq k-1.$ So it suffices to prove that the equality in $\operatorname{PS[q]}(a)_{k}$ holds. Suppose to the contrary that $\operatorname{PS[q]}(a)_{k}$ is strict. For a real number $\delta$, set
$a(\delta):=(a_{1},\cdots,a_{k}-(d+1-k)\delta,a_{k+1}+\delta,\cdots,a_{d+1}+\delta)$
and
$h(\delta):=\frac{f_{d}(a(\delta))}{\Pi_{i=1}^{k-1}a_{i}}=(a_{d+1}+\delta)^{d-k}\cdot(a_{k}-(d+1-k)\delta).$
Arguing similarly as before, we can have $a(\delta)\in X(d,q)$ as $\delta$ varies in a small open neighbourhood of 00. Consider the derivative of $h(\delta)$:
@@
By construction $h(\delta)$ has a local minimal value at $\delta=0$, we must have $h^{\prime}(0)=0$, which implies $a_{d+1}(d+1-k)=a_{k}(d-k)$. Consider the double derivative value of $h(\delta)$ at $\delta=0$:
$h^{\prime\prime}(0)=(d-k)(d-1-k)a_{d+1}^{d-2-k}a_{k}-2(d-k)(d+1-k)a_{d+1}^{d-1-k}.$
@@
for each $l\in\{1,\cdots,d+1\}$. Note that by definition, we have
$y[q](1)$ ${}=(\frac{1}{d+1},\cdots,\frac{1}{d+1}),$
$y[q](d)$ ${}=(\frac{q}{1+u_{1,q}},\cdots,\frac{q}{1+u_{d-1,q}},\frac{q}{2u_{d,q}},\frac{q}{2u_{d,q}}),$
$y[q](d+1)$ ${}=(\frac{q}{1+u_{1,q}},\cdots,\frac{q}{1+u_{d,q}},\frac{q}{u_{d+1,q}}).$
It is easy to see that for each $l\in\{1,\cdots,d+1\}$, $y[q](l)\in X(d,q)$.
@@
Fix two integers $d\geq 2$ and $q\geq 1$. Let $S$ be a $d$-dimensional $\frac{1}{q}$-lc Fano simplex. Suppose $\beta_{1}\geq\beta_{2}\geq...\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}\in\operatorname{int}(S)$ with respect to $S$. Let $f_{d+1}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d+1}(x)=\Pi_{i=1}^{d+1}x_{i}$. Then
$\Pi_{i=1}^{d+1}\beta_{i}\geq f_{d+1}(y[q](d+1))=\frac{q^{d+2}}{u_{d+1,q}^{2}},$
and the equality holds if and only if
$(\beta_{1},\cdots,\beta_{d+1})=(\frac{q}{1+u_{1,q}},\cdots,\frac{q}{1+u_{d,q}},\frac{q}{u_{d+1,q}}).$
@@
One can check directly that (3.6) holds. So we may assume that $d\geq 3$.
In case $q=1$, (3.6) follows from [3, Theorem 4.3 $(c)$]. So in the next discussion, we always assume that $q\geq 2$.
@@
Fix two integers $d\geq 2$ and $q\geq 1$. Let $S$ be a $d$-dimensional $\frac{1}{q}$-lc Fano simplex. Suppose $\beta_{1}\geq\beta_{2}\geq...\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}\in\operatorname{int}(S)$ with respect to $S$. Let $f_{d}:\mathbb{R}^{d+1}\longrightarrow\mathbb{R}$ be the continuous function defined by sending $x=(x_{1},\cdots,x_{d+1})$ to $f_{d}(x)=\Pi_{i=1}^{d}x_{i}$. Then
@@
except for the case $d=2,q=1$. In the exceptional case $d=2,q=1$, $\Pi_{i=1}^{d}\beta_{i}\geq f_{2}(y[q](1))=\frac{1}{9}.$ Furthermore, for $d=3,q=1$ the equality holds if and only if $(\beta_{1},\beta_{2},\beta_{3},\beta_{4})=(\frac{1}{2},\frac{1}{6},\frac{1}{6},\frac{1}{6})$ or $(\frac{1}{2},\frac{1}{3},\frac{1}{12},\frac{1}{12})$; for $d=3,q\geq 2$ or $d\geq 4$, the equality holds if and only if